can you help me solve this equation?

I am trying to solve a system of linear quations by the substitution method.
6x - 4y = 3
3x - 2y = 9
I need to see how you worked the solution so that I can make sure that I am doing it right and can do the rest of the problems on my own.
Thank you!!!

Answer by xcentaur(357)

when you multiply the second equation by 2, we get-
6x - 4y = 3....[1]
3x - 2y = 9 ...[2]


6x - 4y = 3
6x - 4y = 18


Clearly,this is not possible,as both x and y are getting eliminated.Please check your question,and post back again.
Prabhat.